At September 1, 2017, Crews Co.  reported owner’s equity of $136,000. During the month, Crews generated revenues...

  1. At September 1, 2017, Crews Co.  reported owner’s equity of $136,000. During the month, Crews generated revenues of $20,000, incurred expenses of $12,000, purchased equipment for $5,000 and withdrew cash of $2,000. What is the amount of owner’s equity at September 30, 2017?
    1. $136,000
    2. $8,000
    3. $137,000
    4. $142,000
  2. A basic assumption of accounting that requires activities of an entity be kept separate from the activities of its owner is referred to as the
    1. stand alone concept.
    2. monetary unit assumption.
    3. corporate form of ownership.
    4. economic entity assumption
  3. A business pays weekly salaries of $20,000 on Friday for a five-day week ending on that day.  The adjusting entry necessary at the end of the fiscal period ending on a Thursday is
    1. debit Salaries Payable, $16,000; credit Cash, $16,000.
    2. debit Salaries Expense, $16,000; credit Cash, $16,000.
    3. debit Salaries Expense, $16,000; credit Salaries Payable, $16,000.
    4. debit Salaries Expense, $4,000; credit Salaries Payable, $4,000.
  4. RAS Corporation issued a one-year, 12%, $200,000 note on August 31, 2017. Interest expense for the year ended December 31, 2017 was
    1. $24,000.
    2. $10,000.
    3. $8,000.
    4. $6,000.
Answer #1

solution 1:

amount of owner equity on Sept 30, 2017 = Beginning owner equity + Net income - Withdrawls

= $136,000 + ($20,000 - $12,000) - $2,000 = $142,000

Hence option d is correct.

Solution 2:

A basic assumption of accounting that requires activities of an entity be kept separate from the activities of its owner is referred to as the "economic entity assumption"

Hence option d is correct.

Solution 3:

Salary expense accured on thrusday = $20,000*4/5 = $16,000

The adjusting entry necessary at the end of the fiscal period ending on a Thursday is "debit Salaries Expense, $16,000; credit Salaries Payable, $16,000"

Hence option c is correct.

Solution 4:

Interest expense for the year ended December 31, 2017 was = $200,000*12%*4/12 = $8,000

Hence option c is correct.